OpenStudy (anonymous):

Verify that the Intermediate Value theorem applies to the indicated interval and find the value of c guaranteed by the theorem. f(x) = x^2 - 6x + 8, [0,3], f(c) = 0

OpenStudy (anonymous):

@Neemo @jim_thompson5910

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

what is your indicated interval?

OpenStudy (anonymous):

it is 0,3

OpenStudy (anonymous):

i then found that it takes all values between -1 and 8. and 0 iis between these values... so what wwould i put as my answer?

OpenStudy (anonymous):

f(0) = 8 f(3) = -1 f(2) = 0 f(2) belongs to [f(0), f(3)]

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

is that all I put for my answer

jimthompson5910 (jim_thompson5910):

Basically because f(0) = 8 and f(3) = -1, this means that f(c) = 0 must lie between the two somewhere by the intermediate value theorem

OpenStudy (anonymous):

ok so thats basically my answer then, thats it.

jimthompson5910 (jim_thompson5910):

So you can either factor, use a graph, or the quadratic formula to find that f(2) = 0, so c = 2

OpenStudy (anonymous):

ok thx. i have one more, its easier. can you help with this? :

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Describe the difference between a discontinuity that is removable and one that is a nonremovable.? In your explanation, give examples of the following: a) a function with a nonremovable discontinuity at x=2. b) a function with a removable discontinuity at x= -2. c) a function that has both of the characteristics described in a. and b.

jimthompson5910 (jim_thompson5910):

x/(x-2) has a non-removable discontinuity at x = 2 since plugging in x = 2 results in a division by zero error (and you can't simplify to make the error go away)

jimthompson5910 (jim_thompson5910):

something like (3x+6)/(x+2) has a removable discontinuity at x = -2 since (3x+6)/(x+2) (3(x+2))/(x+2) 3 So (3x+6)/(x+2) = 3, but we must restrict that \(x \neq -2\) to make sure that the two domains match up

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

a function that has both x = 2 as a non-removable discontinuity and x = -2 as a removable discontinuity is something like (5x+10)/(x^2-4) since (5x+10)/(x^2-4) (5(x+2))/((x+2)(x-2)) 5/(x-2) So (5x+10)/(x^2-4) = 5/(x-2), but \(x \neq 2\) and \(x \neq -2\) to avoid division by zero errors

OpenStudy (anonymous):

thx so much

jimthompson5910 (jim_thompson5910):

np